OpenStudy (anonymous):

can anybody tell me if 3.9,3.99,3.999,......... is a sequence or not please its really important

OpenStudy (anonymous):

Well, if r = 0.1, you could identify \[a^{n+1}=a^n+r^n*0.9=a^n+0.1^n*0.9\] with \[a_0=3\] right?

OpenStudy (anonymous):

that means its a sequence isnt it

OpenStudy (anonymous):

I gave the answer as you seemed to be in a hurry :)

OpenStudy (anonymous):

I'm afraid I dont know the exact definition of a sequence, could anyone confirm?

OpenStudy (anonymous):

its ok then

OpenStudy (anonymous):

I should have made the sequence number a subscript for clarity: \[a_{n+1}=a_n+r^n*0.9\]

OpenStudy (anonymous):

Well the math checks out as far as I can see. At least with the parameters mentioned. That formula would turn out some weird numbers(at least weird compared to the stated ones) if you went left on the number line xP

OpenStudy (anonymous):

thanks TimSmit

OpenStudy (anonymous):

Hm that is true, only for positive numbers this :) But when you take \[n \in \mathbb{N}\] its ok right?

OpenStudy (anonymous):

yes thats what I wanted to know thanks for helping me

OpenStudy (anonymous):

or to be more clear \[n \in \mathbb{N}_0\]

OpenStudy (anonymous):

No problem, I hope I didn't give the answer too soon.

OpenStudy (anonymous):

Tim sorry sorry the formula you gave me isnt giving me the sequence which I asked for

OpenStudy (mandre):

I know it's not that important but shouldn't \[a _{n+1}=a _{n}+r ^{n}∗0.9\] be \[a _{n+1}=a _{n}+r ^{n}∗9\]

OpenStudy (anonymous):

Yes. I was substituting that automatically for some reason so I didn't even notice it.

OpenStudy (mandre):

Unless \[a _{0}\] should be 3.9 then 0.9 is correct?

OpenStudy (anonymous):

I start counting from 0 most of the times :) Matter of preference I guess...You are right if starting from 1.

OpenStudy (mandre):

Yes. a0 is 3.9 according to mahendra

OpenStudy (anonymous):

Using a1=3.9 and starting at 1 it also checks out I think: \[a_{n+1}=a_n + 0.1^n*0.9\] \[3.99=3.9 + 0.1^1*0.9\] \[3.999=3.99 + 0.1^2*0.9\] etc

OpenStudy (mandre):

my concern is whether there are restrictions on r's value.

OpenStudy (mandre):

Can r be an "equation" instead of just a "number"?

OpenStudy (anonymous):

this is the final answer to my question and Tim you were asking what is sequence any series whose next value can be derived from previous one is called a sequence here n is the only variable and its value only belongs to the set of natural no.

OpenStudy (anonymous):

r is just a constant

OpenStudy (mandre):

So you're saying r should be a constant and your list of numbers is not a sequence?

OpenStudy (anonymous):

no with r as a constant the list of no. we get is a sequence because with r as a constant every next term of the list can be derived from the previous one which is according to the definition of sequence

OpenStudy (mandre):

but r changes so r is not a constant...

OpenStudy (anonymous):

the value of r is not changing its value is constant equal to 0.1 value of r to the power n is changing because n is variable

OpenStudy (anonymous):

It is a sequence. Note that we can re-write the terms as:\[\bf \left\{ 3.9.3.99,3.999.. \right\}=\left\{ 3.9,3.9+0.09,3.99+0.009.. \right\}\]\[\bf = \left\{ 3.9+0.09\left( \frac{ 1 }{ 10 } \right)^0,3.99+0.09\left( \frac{ 1 }{ 10 } \right)^1,3.999+0.09\left( \frac{ 1 }{ 10 } \right)^2.. \right\}\]Hence the sequence is defined as:\[\bf a_n=a_{n-1}+0.09\left( \frac{ 1 }{ 10 } \right)^{n-1}\]Given that \(\bf a_1=3.9+0.09(1/10)^0=3.99\). @m4mahendra20

OpenStudy (anonymous):

Sorry, let me make a correction: I mean to say; "Given that \(\bf a_1=3.9+0.09(1/10)^0=3.9\)"* I accidentally wrote 3.99 in last statement...

OpenStudy (mandre):

So it's an arithmetic sequence?

OpenStudy (anonymous):

i cant really say that the only thing i can say that it is a sequence but perhaps it is not arithmetic

OpenStudy (mandre):

You are adding someting to each preceding term so one would think it would be arithmetic instead of geometric. The only problem is that d (difference, we used r) is not a constant and I don't think you can do that for an arithmetic sequence. \[a _{n}=a+(n-1)d\] but in your case d changes.